OpenStudy (anonymous):

I really need help with this!!!!!! PLEASEEEE Which expresses the given inequality in interval notation? x < 2 A. x E (-infiniti, 2) B. x E (2, infiniti) C. x E [2, infiniti) D. x E (-infiniti, 2]

OpenStudy (freckles):

x less than 2 what does that mean what numbers does that mean?

OpenStudy (anonymous):

] means include the number 2] it means include the number 2 however the notation less so 2 is not include

OpenStudy (sidsiddhartha):

before u do that fu need to know what is\[(~~)=?\\ [~~ ]=?\]

OpenStudy (anonymous):

any number less than 2 is include -1,-1000000, - infinite

OpenStudy (sidsiddhartha):

let theres two numbers a,and b are in the intervals then\[(a,b)=open ~interval=a<x<b\\ [a,b]=closed ~interval=a \le x \le b\]

OpenStudy (anonymous):

so A it should be the answer

OpenStudy (anonymous):

Thank you!!! :)

OpenStudy (sidsiddhartha):

yes because 2 is > than x and if u draw it in a number line then it will go endlessly to negetive infinity

OpenStudy (sidsiddhartha):

so \[ x \in (- \infty,2)\]